bihggn:

0.007592 kg=_____mg 1kg = 1000g 1g = 1000mg

bihggn:

answer please im slow

iYuko:

7592mg, my friend.

Nnesha:

use the given information first convert kg to g and then g to mg \[\large \rm 0.007592 kg \cdot \frac{1000g}{1kg}\]\[\large \rm 0.007592 \cancel{kg} \cdot \frac{1000g}{1\cancel{kg}}= (0.007592 \cdot 1000)g\] Convert g to mg \[\large\rm (0.007592 \cdot 1000)g \cdot \frac{1000 ~mg}{g}\]\[\large\rm (0.007592 \cdot 1000)\cancel{g} \cdot \frac{1000 ~mg}{\cancel{g}}= (0.007592\cdot1000\cdot1000) mg\]